User name R. Michaels
Log entry time 15:55:16 on July25,2005
Entry number 149424
This entry is a followup to: 149033
keyword=cavity monitors in spectr.DAQ mode
Regarding the (new) cavity monitors and the spectrometer DAQ
readout: The oscillations seen on the scope on Friday are normal
with no beam; the X,Y flops around since one is dividing by zero.
During today's controlled access, J. Musson injected signals into
the cavities. It simulates 25 uA and rather "large" displacements
of +/- 5 mm. Spectr.DAQ run 2274 with upstream cavities at
(X,Y) = (1V, 5V) = (-5 mm, +5 mm) very roughly, and run 2275
with (X,Y) = (5V, 1V) = (+5 mm, -5 mm) are shown in fig 1 for
raw signals. Note, this 12-bit ADC ranges to 4096, so the signals
are in the dynamic range.
The next step is to complete the commissioning plan.